Use ELIMINATION to solve each SYSTEM!?
1.) 2x+y=41; 2x+5y=13
____________________
2.) -5x-2y=44; -3x +3y=39

Answer by ankor@dixie-net.com(22740)

1.)
2x + y = 41
2x +5y = 13
-------------Subtraction eliminates x, find y
0 - 4y = 28
divide both sides by -4
y = -7
Use the 1st equation to find x
2x - 7 = 41
2x = 41 + 7
2x = 48
x = 48/2
x = 24
Check solution in the 2nd equation
2(24) + 5(-7) = 13
48 - 35 = 13
:
____________________
2.)
-5x - 2y = 44
-3x + 3y = 39
Multiply the 1st equation by 3, the 2nd equation by 2
-15x - 6y = 132
-6x + 6y = 78
---------------Adding eliminates, find x
-21x = 210
Divide both sides by -21
x = -10
use the 1st equation to find y
-5(-10) - 2y = 44
+50 - 2y = 44
-2y = 44 - 50
-2y = -6
divide both sides by -2
y = +3
Check solutions in the 2nd equation
-3(-10) + 3(3) = 39
30 + 9 = 39